Find the x intercepts for y = x^2 + 3x -10
Answer by mananth(16949)

y = x^2 + 3x -10
for x intercept y=0
x^2+3x-10=0
(x+5)(x-2)=0
x=-5 , x=2
...
m.ananth@hotmail.ca